Personal Electronix sells iPads and iPods. The business is divided into two divisions along product lines. CVP income statements for a recent quarter’s activity are presented below.

iPad Division

iPod Division

Total

Sales $968,500 $521,500 $1,490,000
Variable costs 687,635 286,825 974,460
Contribution margin $280,865 $234,675 515,540
Fixed costs 142,206
Net income $373,334

Determine sales mix percentage and contribution margin ratio for each division. (Round answers to 0 decimal places, e.g. 15%.)

Sales Mix Percentage

iPad division

enter a percentage number rounded to 0 decimal places %

iPod division

enter a percentage number rounded to 0 decimal places %

Contribution Margin Ratio

iPad division

enter a percentage number rounded to 0 decimal places %

iPod division

enter a percentage number rounded to 0 decimal places %

eTextbook and Media

  

  

Calculate the company’s weighted-average contribution margin ratio. (Round computations and final answer to 2 decimal places, e.g. 15.26%.)

Weighted-average contribution margin ratio

enter a percentage number of the weighted-average contribution margin ratio rounded to 2 decimal places %

eTextbook and Media

  

  

Calculate the company’s break-even point in dollars. (Round computations to 2 decimal places and final answer to 0 decimal places, e.g. 1,526.)

Break-even point

$enter the break-even point in dollars rounded to 2 decimal places

eTextbook and Media

  

  

Determine the sales level in dollars for each division at the break-even point. (Round computations to 2 decimal places and final answers to 0 decimal places, e.g. 1,526.)

Break-even point

iPad division

$enter a dollar amount rounded to 2 decimal places

iPod division

$enter a dollar amount rounded to 2 decimal places
Answer #1

Answer- Sales mix- iPad Division = ($968500/$1490000)*100 = 65%

iPod Division = ($521500/$1490000)*100 =35%

Contribution margin ratio- (Contribution margin ratio/Sales)*100

iPad Division = ($280865/$968500)*100 = 29%

iPod Division = ($234675/$521500)*100 = 45%

Answer- Weighted average contribution margin ratio =Contribution margin ratio* Sales mix percentage

= ($29%*65%)+(45%*35%)

= 18.85%+ 15.75%

= 34.6%

Answer- Break-even point in dollars = Fixed costs/ Weighted average contribution margin ratio

= $142206/34.6%

= $411000

Answer- - Break-even point in dollars –Break-even point in dollars* Sales mix

iPad Division = $411000*65% =$267150

iPod Division = $411000*35%= $143850